OpenStudy (anonymous):

What is the remainder when (x3 − 3x2 − 13x + 78) is divided by (x + 4)?

OpenStudy (anonymous):

If you know synthetic division, you can just use that to figure out the remainder. However, if you don't you can plug in x = -4 to find the remainder.

OpenStudy (anonymous):

So, either  or you can evaluate \((-4)^3-3(-4)^2-13(-4)+78.\)

OpenStudy (anonymous):

ok give me a sec

OpenStudy (anonymous):

o ok . o its 15

OpenStudy (anonymous):

so*

OpenStudy (anonymous):

Not quite. I think you made an arithmetic mistake.

OpenStudy (anonymous):

\[(-4)^3-3(-4)^2-13(-4)+78=-64-48+52+78=4+14=18\]

OpenStudy (anonymous):

oooooo

OpenStudy (anonymous):

Also, if you were to do the synthetic method (which I think is easier for these bigger numbers, you'd get: